Taylor Swift wants to know the proportion of her fans who listened to her new single ME! within the first hour of it being released.  In order to estimate this, she takes a sample of 150 of her fans and asks them if they listened to the song in this time period.  Of the 150 fans, 135 of them (90%) responded that they did listen to the song during this time period.  What is the parameter and what is its value?
Answer #1

Answer:

The parameter in this scenario is given as the population proportion of her fans who listened to her new single ME!within the first hour of it being released.  

The estimate for population proportion = p = x/n = 135/150 = 0.9

The estimate for the population proportion is 0.90 or 90%.